connect FINANCE z Connect Quiz4 instructions i help Question 9 (of 20) Dyrdek Enterprises has equity with a market value of $12.5 million and the market value of debt is $4 40 million. The company is evaluating a new project that has more risk than the firm. As a resut the company will apply & fisk od ustment factor of 18 percent The new project will cost $2.54 million today and provide annual cash flows of $661,000 for the next 6 years The companys cost of equity is 1175 percent and the pretax cost of debt is 505 percent The tax rate is 39 percent What is the projects NPV? O s201190 $506.832 O $347.611 5168.636 O S196992 Check my work 818 PM O Type here to search ?
Answer #1

Market Value of equity = E = $12.5 million

Market Value of debt = D = $4.40 million

Weight of Equity = We = E/(E+D) = 12.5 / (12.5 + 4.40) = 0.7396

Cost of equity = Ke = 11.75% = 0.1175

Weight of Debt = Wd = D/(E+D) = 4.40 / (12.5 + 4.40) = 0.2604

Cost of Debt = Kd = 5.05% = 0.0505

Tax rate = T = 39% = 0.39

Weighted average cost of capital = WACC = (We x Ke) + [Wd x (1-T) x Kd ]

WACC = (0.7396 x 0.1175) + [ 0.2604 x (1-0.39) x 0.0505 ] = 0.0949

Given that, an additional adjustment factor of 1.8% is to be added for increased risk.

New WACC = 0.0949 + 0.18 = 0.2749 = r

Now the following steps will follow in order to calculate the Net Present Value:

1. Write the cash flows for years 0-6

2. Compute discount factor for each cash flow. This done by the formula 1/(1+r)n where r is the rate of interest and n is the year.

3. Calculate the Discounted cash flow by finding the product of cash flow and its corresponding discount rate.

4. Add all the discounted cash flows to calculate Net Present Value = $1,844,533.98.

The table below shows the value of all heads described in bold above.

Discount Discounted Year Cashlow factor factor cash flow 2540000 661000 0.7 661000 0.615245 406676.63 661000 661000 0.378526